OpenStudy (anonymous):

help please. Find (f + g)(x), (f - g)(x), (f . g)(x), and (f/g)(x) for each f(x) and g(x). 1. f(x) = a^2 - 1 , g(x) = a/(a + 1) (f+g)(x) (f-g)(x) and (f/g)(x)

OpenStudy (freckles):

(f+g)(x) means to add f(x) and g(x)

OpenStudy (freckles):

(f-g)(x) means to take away g(x) from f(x) or just write f(x)-g(x)

OpenStudy (freckles):

(f/g)(x) means to write f(x)/g(x)

OpenStudy (anonymous):

ok i have a question

OpenStudy (anonymous):

what is wrong with this picture \[f(x) = a^2 - 1\]

OpenStudy (anonymous):

im lost

OpenStudy (freckles):

it is a constant function satellite :p example: \[f(x)=2 \\ g(x)=3+x \\ \text{ then } (f+g)(x)=2+(3+x)=5+x \\ \text{ also } (f-g)(x)=2-(3+x)=-x-1 \\ \text{ and } (\frac{f}{g})(x)=\frac{2}{3+x}\]

OpenStudy (anonymous):

me too if you have a function of x, it should have an x in it

OpenStudy (anonymous):

whats going on?

OpenStudy (freckles):

well are your functions really f(x)=a^2-1 and g(x)=a/(a+1) ?

OpenStudy (anonymous):

don't know does it really say \[f(x)=a^2-1\] or is it \[f(x)=x^2-1\]

OpenStudy (anonymous):

freckles, yes.

OpenStudy (anonymous):

what it says is what it is. I'm just not understanding how to solve it.

OpenStudy (freckles):

whatever they are just add f(x) and g(x) to find (f+g)(x)

OpenStudy (anonymous):

okay

OpenStudy (freckles):

have you found (f+g)(x)?

OpenStudy (anonymous):

nope I've been trying but i don't know what I'm doing

OpenStudy (freckles):

do you know how to add f(x) and g(x) that is do you know how to add a^2-1 and a/(a-1) ?

OpenStudy (anonymous):

no

OpenStudy (freckles):

\[(f+g)(x)=f(x)+g(x) \\ =a^2-1+\frac{a}{a-1}\] that is it unless you are asked to combine the fractions

OpenStudy (anonymous):

no i wasn't.

OpenStudy (freckles):

try this one: \[(f-g)(x)=f(x)-g(x)\]

OpenStudy (anonymous):

no clue

OpenStudy (freckles):

f(x)=a^2-1 and g(x)=a/(a-1) all you have to do is replace f(x) with a^2-1 and g(x) with a/(a-1)

OpenStudy (anonymous):

what exactly am i doing with this? and should the answer be a equation?

OpenStudy (anonymous):

i understand none of this and it gets frustrating

OpenStudy (freckles):

the expression should be an expression of x in this case the expression of x is a constant expression since the expression is really an expression in terms of a

OpenStudy (anonymous):

I'm attempting to follow

OpenStudy (freckles):

pretend f(x)=5 and g(x)=6 (f+g)(x)=f(x)+g(x)=5+6=11 you are just replacing f(x) with 5 and g(x) with 6 (f-g)(x)=f(x)-g(x)=5-6=-1 I just replaced f(x) with 5 and g(x) with 6 (f/g)(x)=f(x)/g(x)=5/6 again I just replaced f(x) with 5 and g(x) with 6

OpenStudy (freckles):

(f*g)(x)=f(x)*g(x)=5*6=30 again I just replaced f(x) with 5 since in my example I let f(x)=5 and I just replaced g(x) with 6 since in my example I let g(x)=6

OpenStudy (anonymous):

ok

OpenStudy (freckles):

ok now you try finding (f-g)(x) for your f(x) and g(x)

OpenStudy (freckles):

remember (f-g)(x) is the same as f(x)-g(x)

OpenStudy (freckles):

just replace f(x) with a^2-1 and g(x) with a/(a-1)